OpenStudy (anonymous):

what is the equation of the tangent line to the graph of f(x)= x^3+1, at the point (1,2)?

OpenStudy (anonymous):

(y-2)={d/dx (x^3 +1) at (1,2)} (x-1) => y-2 = 3(x-1)

OpenStudy (anonymous):

f'(x) = 3x^2 ==> slope = 3 ===> y=3x-1

OpenStudy (anonymous):

can you use the limit process?

OpenStudy (anonymous):

yes you can.... f '(x=)lim(h->0)={[(x+h)^3 +1]-[(x)^3+1]}/h f '(x) = 3x^2 slope=3 y-2 = 3(x-1) y=3x-1